OpenStudy (anonymous):

Determine the radius of a circle whose arc length PLEASE HELP! Measures 57.1 meters and central angle measures 1.46 radians.

OpenStudy (jdoe0001):

\( \bf arc = r\theta \implies r = \cfrac{arc}{\theta}\)

OpenStudy (anonymous):

What does the r represent? and what does the θ represent?

OpenStudy (jdoe0001):

surprisingly enough, r = radius :)

OpenStudy (jdoe0001):

\(\theta\) = angle in radians

OpenStudy (anonymous):

I got 39.1 is that correct?

OpenStudy (jdoe0001):

yes, r = 39.1

OpenStudy (anonymous):

Okay thank you so much you have no Idea how long I've been trying to solve this!

OpenStudy (jdoe0001):

yw